OpenStudy (anonymous):

determine the vertical asymptotes 2x^2+3x+1/((x^2-2)(5-x^2))

OpenStudy (anonymous):

\[\frac{2x^{2}+3x+1}{(x ^{2}-2)\times (5-x ^{2})}\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im not sure if i should set the demonator to 0 or expand it.

OpenStudy (anonymous):

\[\frac{(x+1)(2x+1)}{(x ^{2}-2)(-x ^{2}+5)}\] can't simpfy

OpenStudy (anonymous):

the vertical asymptotes is in the denominator right? how can i solve that?